I need help with figuring out how to work these types of problems. Step by step. For example, here are two examples of problems I need help understanding how to work.
Solve - -2(-1-3y)=7y
solve 7(6=w)=6(2+w)
I know the goal is to get the variable by itself but I am confused. Please advise. Thanks!

Answer by stanbon(75887)

Solve
-2(-1-3y) = 7y
2+6y = 7y
y = 2
-------

solve
7(6+w) = 6(2+w)
42+7w = 12+6w
w = -30
==============
Cheers,
Stan H.
==============